### 15 Mar 2026 v1

# Deviation Tests for a High-dimensional Mean

Abstract
This paper investigates testing for deviation of a high-dimensional mean vector $\boldsymbol{\mu}$. In contrast to the standard one-sample significance test of the form: $H_{0}^{\texttt{e}}:\boldsymbol{\mu}=\boldsymbol{\mu}_{0}$ versus $H_{1}^{\texttt{e}}:\boldsymbol{\mu}\neq\boldsymbol{\mu}_{0}$, we focus on testing the deviation $H_{0}:\|\boldsymbol{\mu}-\boldsymbol{\mu}_{0}\|_{2}\geq d_{0}$ versus $H_{1}:\|\boldsymbol{\mu}-\boldsymbol{\mu}_{0}\|_{2}<d_{0}$ for a prespecified length $d_{0}>0$. Constructing a valid test statistic for this problem is technically nontrivial. By applying the concept of positive and negative feedback processes from control theory, we propose a test statistic based on a two-armed bandit (TAB) process. The deviation test is also extended to the two-sample setting. Simulation experiments confirm a good performance of the tests in finite samples. Finally, a real data analysis demonstrates the practical significance of the proposed deviation tests.
Keywords: High-dimensional significance test; High-dimensional test for deviation of mean; Two-armed bandit (TAB) process.
1 Introduction
The significance test for the hypothesis that a population mean equals a given vector, or testing the equality of means from two populations, has been comprehensively studied in high-dimensional settings. For background, readers may refer to [1, 4, 3, 20, 14] and references therein. Specifically, given a random sample $\{\boldsymbol{x}_{t}\in\mathbb{R}^{n}:t=1,\dots,T\}$ with population mean vector $\boldsymbol{\mu}$, where the data dimension $n$ and sample size $T$ tend to infinity proportionally, the standard significance test is:
$H_{0}^{\texttt{e}}:\boldsymbol{\mu}=\boldsymbol{\mu}_{0}\quad\text{versus}\quad H_{1}^{\texttt{e}}:\boldsymbol{\mu}\neq\boldsymbol{\mu}_{0},$ (1)
where $\boldsymbol{\mu}_{0}\in\mathbb{R}^{n}$ is a given reference mean vector. However, this equality test is often too restrictive for real-world applications. First, the case where $\boldsymbol{\mu}=\boldsymbol{\mu}_{0}$ is rare in practice, especially in high dimensions. Many empirical studies (e.g., [13, 14, 9]) find strong evidence to reject $H_{0}^{\texttt{e}}$.
Moreover, rejecting $H_{0}^{\texttt{e}}$ does not provide information regarding the magnitude of the difference between $\boldsymbol{\mu}$ and $\boldsymbol{\mu}_{0}$. Such information is however critical in biopharmaceutical statistical analysis. For example, when comparing a medicine to its biosimilar, researchers aim to prove the biosimilar is clinically “equivalent” to the original. Let $\boldsymbol{\mu}_{0}\in\mathbb{R}^{n}$ be a vector representing the clinical assessments of the original medicine, and $\{\boldsymbol{x}_{t}\in\mathbb{R}^{n}:t=1,\dots,T\}$ be the assessments of the biosimilar from $T$ independent clinical trials. We treat $\boldsymbol{x}_{t}$ as random vectors with population mean $\boldsymbol{\mu}\in\mathbb{R}^{n}$. To verify the efficacy of the biosimilar, we expect $\boldsymbol{\mu}$ to be close to the benchmark $\boldsymbol{\mu}_{0}$ (e.g., $\|\boldsymbol{\mu}-\boldsymbol{\mu}_{0}\|_{2}<d_{0}$ for a constant $d_{0}>0$). Therefore, we consider the following deviation test:
$H_{0}:\|\boldsymbol{\mu}-\boldsymbol{\mu}_{0}\|_{2}>d_{0}\quad\text{versus}\quad H_{1}:\|\boldsymbol{\mu}-\boldsymbol{\mu}_{0}\|_{2}\leq d_{0}.$ (2)
Compared to the equality test, (2) provides more concrete information about the relationship between $\boldsymbol{\mu}$ and $\boldsymbol{\mu}_{0}$, making it more appropriate for testing similarity. Rejecting $H_{0}$ implies that the biosimilar is clinically similar to the original medicine (with a difference smaller than $d_{0}$ in $L_{2}$ norm), whereas accepting $H_{0}$ indicates they are not clinically equivalent.
However, even when $\{\boldsymbol{x}_{t}\}$ are univariate random variables, very few studies (e.g., [6]) have examined the test (2) because constructing an appropriate test statistic is technically difficult. For instance, if $\{\boldsymbol{x}_{t}\}$ are independent and $\mathcal{N}(\mu,\sigma^{2})$—distributed (i.i.d.), the test simplifies to $H_{0}:|\mu-\mu_{0}|>d_{0}$ versus $H_{1}:|\mu-\mu_{0}|\leq d_{0}$. Since the parameter spaces under $H_{0}$ and $H_{1}$ share the same dimension, the classic likelihood ratio test cannot be applied in the standard way.
Notably, utilizing the idea of positive/negative feedback processes from control theory, [6] proposed a novel sequential framework for constructing a test statistic for (2) in the univariate case based on a two-armed bandit (TAB) process. Given a sequence of univariate random samples $\{x_{t}\in\mathbb{R}:t=1,\dots,T\}$, the test statistic $\{\hat{b}_{T,t}:t=1,\dots,T\}$ is constructed sequentially:
$\hat{b}_{T,t}=\left\{\begin{array}[]{ll}h(x_{1}),&t=1,\\ \hat{b}_{T,t-1}+h(x_{t}),&t=2,\dots,T,\end{array}\right$. (5)
where $h(\cdot)$ is a well-chosen innovation function. This construction is similar to the univariate sequential probability ratio test (SPRT) for the sample mean [17, 18]:
$H_{0}^{\texttt{SPRT}}:\mu=\mu_{0}\quad\text{versus}\quad H_{1}^{\texttt{SPRT}}:\mu=\mu_{1},\quad\text{for }\mu_{0}\neq\mu_{1}.$
Recent developments in SPRT can be found in [7, 11, 10]. In SPRT, the key is constructing a confidence interval to control Type I/Type II errors for a given significance level $\alpha$, while the innovation function $h(\cdot)$ is usually simple, such as the log-likelihood function.
In contrast, [6] designed the innovation function $h(\cdot)$ to exhibit specific properties under $H_{0}$ and $H_{1}$. Precisely, the test statistic $\hat{b}_{T,t}$ follows a negative feedback process under $H_{0}$ (asymptotically concentrating around 0) and a positive feedback process under $H_{1}$ (asymptotically diverging to infinity). In short, the innovation function $h(\cdot)$ drives distinct asymptotic behaviors for $\hat{b}_{T,t}$ under the null and alternative hypotheses, which is the key for solving (2) in the univariate case.
In this article, we propose a systematic procedure to construct a statistic for the deviation test (2) in high-dimensional settings. Due to the “curse of dimensionality,” directly extending hypothesis tests from one-dimensional to high-dimensional cases is difficult. While one might consider examining (2) using multiple testing literature—where procedures like [2] control the False Discovery Rate (FDR) for test (1) (see [16, 19, 22, 8])—adapting these procedures to test (2) is not straightforward. Particularly, the non-null distance $d_{0}$ in (2) is a global distance that cannot be easily distributed across $n$ dimensions.
Technically, while we use the same sequential framework as [6], constructing a proper innovation function $h(\cdot)$ in high dimensions presents new challenges. Notably in the univariate case, the statistic $\hat{b}_{T,t}$ in (5) is independent of the innovation $h(x_{t})$ at each step $t$. However in high dimensions, in order to ensure the statistic performs as a negative/positive feedback process under $H_{0}/H_{1}$, our proposed innovation function relies on both $\boldsymbol{x}_{t}$ and prior samples $\boldsymbol{x}_{s}$ (where $s<t$). This creates a high correlation between the statistic and the innovation, significantly complicating a theoretical analysis.
The remainder of this paper is organized as follows. Section 2 studies the high-dimensional deviation test (2) for the one-sample case. Section 3 extend the test to the two-sample case. Section 4 presents numerical experiments evaluating the finite-sample performance of the proposed test. Section 5 provides an empirical study. All proofs of our main results are postponed to the Appendix.
We conclude this section by listing useful notations:
  1. 1.
    For two sequences $\{a_{n}\}$ and $\{b_{n}\}$, $a_{n}\leq O(b_{n})$ implies $a_{n}\leq Cb_{n}$ for some constant $C>0$ as $n$ becomes sufficiently large. The equality $a_{n}=O(b_{n})$ means $a_{n}\leq O(b_{n})$ and $b_{n}\leq O(a_{n})$.
  2. 2.
    Given a matrix $\mathbf{A}=[A_{i,j}]_{n\times n}$, $\text{Tr}(\mathbf{A})=\sum_{i=1}^{n}A_{i,i}$ and $\mathbf{A}^{\prime}$ denotes the transpose. $\|\mathbf{A}\|$ denotes the spectral norm.
  3. 3.
    The norm $\|\cdot\|_{2}$ is the $L_{2}$ norm of vectors.
  4. 4.
[…]
  5. 5.
    For sequences of random variables $\{X_{n}\}$ and $\{Y_{n}\}$, $X_{n}\overset{d}{\sim}Y_{n}$ as $n\to\infty$ means that for any $x\geq 0$, $\lim_{n\to\infty}|\mathbb{P}(|X_{n}|\leq x)-\mathbb{P}(|Y_{n}|\leq x)|=0$.
2 One-sample case
Since the reference mean $\boldsymbol{\mu}_{0}$ is given in practice, without loss of generality, we assume $\boldsymbol{\mu}_{0}=\boldsymbol{0}$ in (2). With $\boldsymbol{\mu}_{0}=\boldsymbol{0}$, the test (2) becomes
$H_{0}:\|\boldsymbol{\mu}\|_{2}>d_{0}\quad\text{versus}\quad H_{1}:\|\boldsymbol{\mu}\|_{2}\leq d_{0}.$ (6)
Moreover, we adopt the following asymptotic regime.
Assumption 1.
Both the sample size $T$ and the dimension $n=n(T)$ of $\boldsymbol{x}_{t}$ grow to infinity such that
$\lim_{T\to\infty}n/T=c_{0}\in[0,\infty).$
Note that our main results are valid under $\lim_{T\to\infty}n/T=0$, which includes the classical low-dimensional case with fixed $n$ and $T\to\infty$.
To construct the statistic for (6), let $T=T_{1}+T_{2}$ such that $\lim_{T\to\infty}T_{1}/T=c_{1}\in(0,1)$. Define
$\left\{\begin{array}[]{l}\displaystyle X_{t}=\frac{1}{T_{1}}\sum_{s=1}^{T_{1}}(\boldsymbol{x}_{s}^{\prime}\boldsymbol{x}_{t+T_{1}}-d_{0}^{2}),\quad t=1,\cdots,T_{2},\\[15.0pt] \displaystyle\hat{\tau}_{1}:=\frac{1}{T_{2}}\sum_{t=1}^{T_{2}}X_{t}\quad\text{and}\quad\hat{\sigma}_{1}^{2}:=\frac{1}{T_{2}}\sum_{t=1}^{T_{2}}X_{t}^{2}-\hat{\tau}_{1}^{2}.\end{array}\right$. (9)
Let
$\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t}):=\frac{1}{T_{2}}\sum_{s=1}^{t}\theta_{s}X_{s}+\frac{1}{\sqrt{T_{2}}}\sum_{s=1}^{t}\frac{\theta_{s}X_{s}}{\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}},\quad t=1,\cdots,T_{2},$ (10)
where $\vec{\theta}_{t}=(\theta_{1},\cdots,\theta_{t})^{\prime}$ such that $\theta_{1}=1$ and for $t=2,\cdots,T_{2}$
$\theta_{t}=\left\{\begin{array}[]{cc}1,&\text{if }\ \mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})\leq 0,\\[6.0pt] -1,&\text{if }\ \mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})>0.\end{array}\right$.
Note that the construction process (10) satisfies that for $t\geq 2$
$\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})=\mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t})+\frac{\theta_{t}X_{t}}{\sqrt{T_{2}}}\left(\frac{1}{\sqrt{T_{2}}}+\frac{1}{\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}}\right):=\mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t})+\tilde{h}(X_{t}),$ (11)
where $\mathbb{E}[X_{t}]=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ by (9). Intuitively, under $H_{0}$, $\mathbb{E}[X_{t}]\geq 0$, and the parameter $\theta_{t}$ makes the innovation $\tilde{h}(X_{t})$ to have in average an opposite sign to the current value $\mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})\leq 0$ and vice versa, which defines a negative feedback process. By contrast, $\{\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})\}$ acts as a positive feedback under $H_{1}$, where $\mathbb{E}[X_{t}]<0$. Consequently, as $t\to\infty$, the negative feedback process will restrict our statistic $\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})$ asymptotically concentrating around 0 under $H_{0}$, while the positive feedback process will push $\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})$ asymptotically diverging to infinity under $H_{1}$.
Actually, $\big\{\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})\big\}$ follows a TAB process and $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ asymptotically converges to the bandit distribution $\mathcal{B}(\kappa)$, whose density function has the form of
$f^{\kappa}(x):=\frac{1}{\sqrt{2\pi}}e^{-(|x|-\kappa)^{2}/2}-\kappa e^{2\kappa|x|}\Phi(-|x|-\kappa),$ (12)
where $\Phi(\cdot)$ is the standard normal cumulative distribution function and $\kappa\in\mathbb{R}$ is a constant such that $\operatorname{sign}(\kappa)=\operatorname{sign}(d_{0}-\|\boldsymbol{\mu}\|_{2})$. Readers can refer to Figure 1 for a density plot of bandit distributions under different values of $\kappa$.
Figure 1: Density plot of bandit distributions under different values of $\kappa$. The continuous curve ($\kappa=0$) is the standard normal distribution.
Notably, when $\kappa=0$, $\mathcal{B}(0)$ is the standard normal. When $\kappa<0$, $\mathcal{B}(\kappa)$ is unimodal and more concentrated around 0 than the standard normal distribution, while when $\kappa>0$, $\mathcal{B}(\kappa)$ has two modes dispersing away from 0. Therefore, the key distinction between the two hypotheses in (2) is the different forms of $\mathcal{B}(\kappa)$’s density under $\kappa<0$ ($H_{0}$) and $\kappa\geq 0$ ($H_{1}$).
We now specify the data generation process.
Assumption 2.
Suppose $\boldsymbol{x}_{t}=\boldsymbol{\mu}+\boldsymbol{\Gamma}\boldsymbol{y}_{t}\in\mathbb{R}^{n}$ such that $\boldsymbol{\Gamma}\boldsymbol{\Gamma}^{\prime}=\boldsymbol{\Sigma}\in\mathbb{R}^{n\times n}$ and $\|\boldsymbol{\mu}\|_{2},\|\boldsymbol{\Sigma}\|\leq O(1)$ and $\liminf_{T\to\infty}\frac{1}{T}\operatorname{Tr}(\boldsymbol{\Sigma})>0$. Moreover, $\boldsymbol{y}_{t}=(y_{1,t},\cdots,y_{m,t})^{\prime}\in\mathbb{R}^{m}$, where all $y_{j,t}$ are independent such that $\mathbb{E}[y_{j,t}]=0,\mathbb{E}[y_{j,t}^{2}]=1$ for $j=1,\cdots,m$ and $t=1,\cdots,T$ and $\kappa_{4}:=\sup_{j,t}\big|\mathbb{E}[y_{j,t}^{4}]-3\big|<\infty$.
Remark 2.1.
We are only interested in the case where $\|\boldsymbol{\mu}\|_{2}\leq O(1)$, since if $\lim_{n\to\infty}\|\boldsymbol{\mu}\|_{2}=\infty$, the hypothesis test (2) is trivial. Moreover, for moment conditions on $y_{j,t}$ in Assumption 2, they coincide with those in some well-known reference like [1, 4].
Theorem 2.1.
Under Assumptions 1 and 2, let us with $\tau_{1}:=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ and $\sigma_{1}^{2}:=\boldsymbol{\mu}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\mu}+\frac{1}{T_{1}}\operatorname{Tr}(\boldsymbol{\Sigma}^{2})$, we have as $T_{2}\to\infty$
$\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\overset{d}{\sim}\mathcal{B}(-\kappa_{1,T_{2}}),\quad\kappa_{1,T_{2}}:=\tau_{1}\big(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})}\big).$ (13)
[…]
Next, given a significance level $\alpha\in(0,1)$, let $z_{\alpha/2}$ be the upper $\alpha/2$ quantile of the standard normal, the rejection region of our hypothesis test (6) is
$\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|>z_{\alpha/2}.$
For the size and power of our proposed statistic $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$, Theorem 2.1 implies that
$\lim_{T\to\infty}\big|\mathbb{P}\big(|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})|>z_{\alpha/2}\big)-\mathbb{P}(|\mathcal{B}(-\kappa_{1,T_{2}})|>z_{\alpha/2})\big|=0,$
it suffices to analyze the tail probabilities of the bandit distribution in (12) as follows:
$g(\kappa):=\mathbb{P}(|\mathcal{B}(-\kappa)|>z_{\alpha/2})=1-\Phi(\kappa+z_{\alpha/2})+e^{-2z_{\alpha/2}\kappa}\Phi(\kappa-z_{\alpha/2}).$ (14)
Intuitively, according to Figure 1, $\mathbb{P}(|\mathcal{B}(\kappa)|>z_{\alpha/2})$ is increasing in $\kappa\in\mathbb{R}$, so $g(\kappa)=\mathbb{P}(|\mathcal{B}(-\kappa)|>z_{\alpha/2})$ would be decreasing in $\kappa\in\mathbb{R}$. Actually, by direct calculations,
$g^{\prime}(\kappa)=-2z_{\alpha/2}e^{-2z_{\alpha/2}\kappa}\Phi(\kappa-z_{\alpha/2})<0,$
which demonstrates that $g(\kappa)$ is decreasing in $\kappa\in\mathbb{R}$.
Size of the test: Under $H_{0}$, since $\tau_{1}=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}>0$, then $\kappa_{1,T_{2}}\geq 0$ by (13) and
$\mathbb{P}(|\mathcal{B}(-\kappa_{1,T_{2}})|>z_{\alpha/2})=g(\kappa_{1,T_{2}})\leq g(0)=1-\Phi(z_{\alpha/2})+\Phi(-z_{\alpha/2})=\alpha,$
so the asymptotic size of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ satisfies
$\lim_{T\to\infty}\mathbb{P}\big(\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|>z_{\alpha/2}\big|H_{0}\big)\leq\alpha.$
Power of the test: Under $H_{1}$, we have $\tau_{1}=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}\leq 0$. Note that $\tau_{1}=\tau_{1}(T)$ is a function of the sample size $T$. Let’s analyze the asymptotic power of the test in the following three cases:
  • Case 1. $\liminf_{T\to\infty}\tau_{1}^{2}T=\infty$. Since $\sigma_{1}^{2}=\boldsymbol{\mu}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\mu}+\frac{1}{T_{1}}\operatorname{Tr}(\boldsymbol{\Sigma}^{2})=O(1)$ by Assumption 2, then $\liminf_{T\to\infty}\kappa_{1,T_{2}}=$-$\infty$ by (13). Using (14), the asymptotic power of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ is full, that is
    $\lim_{T\to\infty}\mathbb{P}\big(\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|>z_{\alpha/2}\big|H_{1}\big)=1.$
  • Case 2. $\liminf_{T\to\infty}\tau_{1}^{2}T\in(0,\infty)$. By (13), there exists a constant $C_{0}>0$ such that
    $\liminf_{T\to\infty}|\kappa_{1,T_{2}}|=\liminf_{T\to\infty}\sqrt{\tau_{1}^{2}T/\sigma_{1}^{2}}>C_{0}.$
    Combining with (14), the asymptotic power of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ satisfies
    $\limsup_{T\to\infty}\mathbb{P}\big(\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|>z_{\alpha/2}\big|H_{1}\big)\leq g(C_{0})<1.$
  • Case 3. If $\liminf_{T\to\infty}\tau_{1}^{2}T=0$, we have $\liminf_{T\to\infty}|\kappa_{1,T_{2}}|=0$ and the asymptotic power of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ satisfies
    $\lim_{T\to\infty}\mathbb{P}\big(\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|>z_{\alpha/2}\big|H_{1}\big)=\alpha.$
In summary, the power of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ is asymptotically 1 when $\liminf_{T\to\infty}\tau_{1}^{2}T=\infty$, and otherwise it lies in the interval $(\alpha,1)$. Readers can refer to Figure 2 for an illustration of the asymptotic the power of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ in these three cases.
Figure 2: Density plot of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\overset{d}{\sim}\mathcal{B}(-\kappa_{1,T_{2}})$ under $H_{1}$, where the red area is the rejection region $|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})|>z_{\alpha/2}$. In case 1 as $|\kappa_{1,T_{2}}|\to\infty$, the blue dashed line suggests the densities of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ almost lie in the rejection region, so the asymptotic power is full. In case 2 as $|\kappa_{1,T_{2}}|\to O(1)$, the purple dotted line shows that only part of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$’s distributions lies in the rejection region, so the asymptotic power is in $(\alpha,1)$. In case 3 as $|\kappa_{1,T_{2}}|\to 0$, $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ is asymptotically normal (continuous curve) and the asymptotic power equals $\alpha$.
3 Two-sample case
In this section, we extend the proposed one-sample test in §2 to the two-sample case. Since the basic framework is nearly the same, we only highlight the key differences and omit repetitive details for simplicity.
Consider two independent $n$-dimensional datasets $\{\boldsymbol{x}_{j_{1}}:j_{1}=1,\cdots,M_{1}\}$ and $\{\boldsymbol{z}_{j_{2}}:j_{2}=1,\cdots,M_{2}\}$ such that all $\{\boldsymbol{x}_{j_{1}},\boldsymbol{z}_{j_{2}}\}$ are independent with $\mathbb{E}[\boldsymbol{x}_{j_{1}}]=\boldsymbol{\mu}_{1},\mathbb{E}[\boldsymbol{z}_{j_{2}}]=\boldsymbol{\mu}_{2}$ and $\mathbb{E}[(\boldsymbol{x}_{j_{1}}-\boldsymbol{\mu}_{1})(\boldsymbol{x}_{j_{1}}-\boldsymbol{\mu}_{1})^{\prime}]=\boldsymbol{\Sigma}_{1},\mathbb{E}[(\boldsymbol{z}_{j_{2}}-\boldsymbol{\mu}_{2})(\boldsymbol{z}_{j_{2}}-\boldsymbol{\mu}_{2})^{\prime}]=\boldsymbol{\Sigma}_{2}$ for $j_{1}=1,\cdots,M_{1}$ and $j_{2}=1,\cdots,M_{2}$. Moreover, the sample sizes $M_{1},M_{2}$ satisfy that
$\lim_{M_{1}\to\infty}n/M_{1}=c_{1}\in[0,\infty)\quad\text{and}\quad\lim_{M_{1}\to\infty}M_{1}/M_{2}=c_{2}\in(0,\infty).$ (15)
Next, similar to the previous one-sample case, we split $\{\boldsymbol{x}_{j_{1}}:j_{1}=1,\cdots,M_{1}\}$ and $\{\boldsymbol{z}_{j_{2}}:j_{2}=1,\cdots,M_{2}\}$ into two subsamples, respectively. Precisely, let $N_{0}:=N_{0}(M_{1})\in\mathbb{N}^{+}$ such that $N_{0}<\min\{M_{1},M_{2}\},\lim_{M_{1}\to\infty}N_{0}=\infty$ and
$\lim_{M_{1}\to\infty}(M_{1}-N_{0})/M_{1}=r_{1}\in(0,1)\quad\text{and}\quad\lim_{M_{2}\to\infty}(M_{2}-N_{0})/M_{2}=r_{2}\in(0,1).$ (16)
For any constant $d_{0}>0$, consider testing the hypotheses
$H_{0,\delta}:\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}>d_{0}\quad\text{versus}\quad H_{1,\delta}:\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}\leq d_{0}.$
First, let $m_{i}:=M_{i}-N_{0}$ for $i=1,2$ and
$\Delta_{0}:=\frac{1}{m_{1}}\sum_{j=1}^{m_{1}}\boldsymbol{x}_{j}-\frac{1}{m_{2}}\sum_{j=1}^{m_{2}}\boldsymbol{z}_{j},$
and
$\left\{\begin{array}[]{l}\displaystyle\Delta_{i}:=\boldsymbol{x}_{m_{1}+i}-\boldsymbol{z}_{m_{2}+i},\quad\text{and}\quad Y_{i}:=\Delta_{0}^{\prime}\Delta_{i}-d_{0}^{2}\quad\text{for}\quad i=1,\cdots,N_{0},\\[6.0pt] \displaystyle\hat{\tau}_{1,\delta}:=\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}Y_{i}\quad\text{and}\quad\hat{\sigma}_{1,\delta}^{2}:=\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}Y_{i}^{2}-\hat{\tau}_{1,\delta}^{2}.\end{array}\right$.
Next, define
$\mathcal{M}_{j,N_{0}}^{\delta}(\vec{\theta}_{j})=\frac{1}{N_{0}}\sum_{i=1}^{j}\theta_{i}Y_{i}+\frac{1}{\sqrt{N_{0}}}\sum_{i=1}^{j}\frac{\theta_{i}Y_{i}}{\sqrt{\hat{\tau}_{1,\delta}^{2}+\hat{\sigma}_{1,\delta}^{2}}},\quad j=1,\cdots,N_{0},$ (17)
where $\theta_{1}=1$ and for $j=2,\cdots,N_{0}$
$\theta_{j}=\left\{\begin{array}[]{cl}1,&\text{if }\ \mathcal{M}_{j-1,N_{0}}^{\delta}(\vec{\theta}_{j-1})\leq 0,\\[6.0pt] -1,&\text{if }\ \mathcal{M}_{j-1,N_{0}}^{\delta}(\vec{\theta}_{j-1})>0,\end{array}\right$.
The following data generation process parallels that of one-sample case in Assumption 2.
Assumption 3.
Let $\boldsymbol{x}_{j_{1}}=\boldsymbol{\mu}_{1}+\boldsymbol{\Gamma}_{1}\boldsymbol{y}_{j_{1}}$ and $\boldsymbol{z}_{j_{2}}=\boldsymbol{\mu}_{2}+\boldsymbol{\Gamma}_{2}\boldsymbol{w}_{j_{2}}$ such that for $i=1,2$, $\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}=\boldsymbol{\Sigma}_{i}$ and $\|\boldsymbol{\mu}_{i}\|_{2},\|\boldsymbol{\Sigma}_{i}\|\leq O(1)$ and $\liminf_{M_{i}\to\infty}\frac{1}{M_{i}}\operatorname{Tr}(\boldsymbol{\Sigma}_{i})>0$. Moreover, $\boldsymbol{y}_{j_{1}}=(y_{1,j_{1}},\cdots,y_{m_{1},j_{1}})^{\prime}\in\mathbb{R}^{m_{1}}$, where all $y_{i,j_{1}}$ are independent random variables such that $\mathbb{E}[y_{i,j_{1}}]=0,\mathbb{E}[y_{i,j_{1}}^{2}]=1$ and $\sup_{i,j_{1}}\mathbb{E}[y_{i,j_{1}}^{4}]<\infty$ for $i=1,\cdots,m_{1}$ and $j_{1}=1,\cdots,T$. Similar conditions are also assumed on the vectors $\boldsymbol{w}_{j_{2}}$.
Theorem 3.1.
Under (15), (16) and Assumption 3, let $\tau_{1,\delta}=\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}-d_{0}^{2}$ and
$\sigma_{1,\delta}^{2}=(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2})^{\prime}(\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2})(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2})+\operatorname{Tr}(\{\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}\}\{\boldsymbol{\Sigma}_{1}/m_{1}+\boldsymbol{\Sigma}_{2}/m_{2}\}).$
Then as $N_{0}\to\infty$,
$\mathcal{M}_{N_{0},N_{0}}^{\delta}(\vec{\theta}_{N_{0}})\overset{d}{\sim}\mathcal{B}(-\kappa_{1,\delta}),\quad\kappa_{1,\delta}:=\tau_{1,\delta}\big(1+N_{0}^{1/2}/(\tau_{1,\delta}^{2}+\sigma_{1,\delta}^{2})^{1/2}\big).$
The proof of Theorem 3.1 is the same as the proof of Theorem 2.1, thus omitted. Finally, given a significance level $\alpha\in(0,1)$, we reject $H_{0,\delta}$ if
$\big|\mathcal{M}_{N_{0},N_{0}}^{\delta}(\vec{\theta}_{N_{0}})\big|>z_{\alpha/2}.$
Asymptotic results of this test parallel previous ones in the one-sample case. In particular, and the asymptotic size of $\mathcal{M}_{N_{0},N_{0}}^{\delta}(\vec{\theta}_{N_{0}})$ is upper-bounded by $\alpha$, and under $H_{1,\delta}$, if we suppose $\liminf_{N_{0}\to\infty}\tau_{1,\delta}^{2}N_{0}=\infty$, the asymptotic power of $\mathcal{M}_{N_{0},N_{0}}^{\delta}(\vec{\theta}_{N_{0}})$ is 1.
4 Numerical experiments
We conduct a numerical experiment to verify the finite-sample performance of our test (6) in the one sample case. Here, we set the significance level $\alpha=0.05$ and $\boldsymbol{x}_{t}\overset{i.i.d.}{\sim}\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}=[0.5^{-|i-j|}]_{i,j}\in\mathbb{R}^{n\times n}$ is the population covariance matrix. Moreover, we split the data to two halves with $T_{1}=T_{2}=T/2$. Results for the two-sample case are highly expected to be similar thus not considered.
Let $\boldsymbol{\mu}=(n^{-1/2},\cdots,n^{-1/2})^{\prime}\in\mathbb{R}^{n}$ be the population mean vector with $\|\boldsymbol{\mu}\|_{2}=1$. Under different $(n,T)$ and $d_{0}=0.5+0.1l$ for $l=0,1,\cdots,10$, we compute the empirical values of $\mathbb{P}\big(|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})|>z_{\alpha/2}\big)$ based on 100 independent samples, all results are summarized in Table 1.
Table 1: Empirical values of $\mathbb{P}\big(|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})|\geq z_{\alpha/2}\big)$ under significance level of $\alpha=0.05$. The null hypothesis is $d_{0}<1$” and the alternative hypothesis is $d_{0}\geq 1$”.
$(n,T)$ $d_{0}$ $0.5$ $0.6$ $0.7$ $0.8$ 0.9 1.0 1.1 1.2 1.3 1.4 1.5
$(100,200)$ 0 0 0 0.01 0.05 0.16 0.42 0.71 0.90 0.98 1
$(200,400)$ 0 0 0 0 0.01 0.14 0.52 0.85 0.98 1 1
$(400,800)$ 0 0 0 0 0.01 0.13 0.71 0.96 1 1 1
$(600,1200)$ 0 0 0 0 0 0.09 0.75 0.99 1 1 1
$(100,150)$ 0 0 0.01 0.03 0.06 0.17 0.34 0.59 0.77 0.92 0.98
$(200,300)$ 0 0 0 0 0.02 0.16 0.47 0.8 0.97 1 1
$(400,600)$ 0 0 0 0 0.01 0.15 0.53 0.93 1 1 1
$(600,900)$ 0 0 0 0 0.01 0.09 0.7 0.96 1 1 1
Recall our discussions about the asymptotic size and power of $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ in the end of §2, since
$\lim_{T\to\infty}\big|\mathbb{P}(|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})|>z_{\alpha/2})-\mathbb{P}(|\mathcal{B}(-\kappa_{1,T_{2}})|>z_{\alpha/2})\big|=0,$
when $d_{0}>\|\boldsymbol{\mu}\|_{2}=1$, we have $\kappa_{1,T_{2}}=O(-\sqrt{T})$, so the empirical values of $\mathbb{P}\big(|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})|>z_{\alpha/2}\big)$ increasingly tend to 1 as $(n,T)$ tend to infinity proportionally. On the other hand, when $d_{0}<\|\boldsymbol{\mu}\|_{2}=1$, then $\kappa_{1,T_{2}}=O(\sqrt{T})$ and the empirical values of $\mathbb{P}\big(|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})|>z_{\alpha/2}\big)$ is decreasing and asymptotically smaller than the significance level $\alpha=0.05$ as $(n,T)$ tend to infinity proportionally. These results well confirm the theoretical findings in §2.
5 Empirical study
We exam an intestinal microbiota dataset in [12]. This dataset includes 1006 samples and each sample contains 130 genus-like phylogenetic groups that cover the majority of the known bacterial diversity of the human intestine. Based on this dataset, [21] and [9] have investigated whether the population mean of microbiome compositions differs between the younger age group and the elder age group. Both of these two papers conclude that this difference is significant but provide no information on the magnitude of this difference. Clearly, it is important to quantify this difference of microbiome compositions between the younger group and the elder group. Thus, we reanalyze this dataset based on our two-sample deviation test developed in §3.
Here, we select individuals aged 35 to 50 as the elder group (313 samples in total) and those under 35 as the younger group (303 samples in total), and consider the logarithm of the raw data. Let $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$ be the population mean vectors of the younger and the elder group, respectively. We then consider the following test:
$H_{0}:\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}>d_{0}\quad\text{versus}\quad H_{1}:\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}\leq d_{0}.$
Let $d_{0}=1.4+0.02i$ for $i=0,1,\cdots,10$, we apply our two-sample deviation test in §3, choosing $N_{0}=100$, then compute the statistic $|\mathcal{M}_{N_{0},N_{0}}^{\delta}|$ in (17) and the corresponding p-values $\mathbb{P}\big(|\mathcal{N}(0,1)|>|\mathcal{M}_{N_{0},N_{0}}^{\delta}|\big)$ under different values of $d_{0}$. The obtained values of the test statistic and corresponding p-values are given in Table 2.
Table 2: Values of $|\mathcal{M}_{N_{0},N_{0}}^{\delta}|$ and $\mathbb{P}\big(|\mathcal{N}(0,1)|>|\mathcal{M}_{N_{0},N_{0}}^{\delta}|\big)$ under different $d_{0}$.
$d_{0}$ 1.4 1.42 1.44 1.46 1.48 1.5 1.52 1.54 1.56 1.58 1.6
$\big|\mathcal{M}_{N_{0},N_{0}}^{\delta}\big|$ 1.61 1.67 1.74 1.82 1.91 2 2.07 2.12 2.19 2.27 2.34
p-value 0.11 0.10 0.08 0.07 0.06 0.05 0.04 0.03 0.03 0.02 0.02
If we set the significance level $\alpha=0.05$, then the critical value is $z_{\alpha/2}=1.96$. According to Table 2, when $d_{0}\geq 1.5$, we will reject $\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}>d_{0}$. Following the theory developed in §3, we conclude that $\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}$ is nearly 1.5 at risk level of 0.05.
Appendix A Appendix
We prove Theorem 2.1 via following steps:
  • In §A.1, for variables $\hat{\tau}_{1}$ and $\hat{\sigma}_{1}$ defined in (9), we show that they are consistent estimators of $\tau_{1}$ and $\sigma_{1}$ in Theorem 2.1, respectively.
  • The whole proof of Theorem 2.1 is provided in §A.2.
A.1 Auxiliary lemmas
Lemma A.1.
Under Assumptions 1 and 2, we have
$\mathbb{E}[(\hat{\tau}_{1}-\tau_{1})^{4}]\leq O(T^{-2})\quad\text{and}\quad\mathbb{E}[(\hat{\sigma}_{1}^{2}-\sigma_{1}^{2})^{2}]\leq O(T^{-1}),$
where $\hat{\tau}_{1},\tau_{1}$ and $\hat{\sigma}_{1},\sigma_{1}$ are defined in (9) and Theorem 2.1.
Proof.
Let’s first show that $\mathbb{E}[(\hat{\tau}_{1}-\tau_{1})^{4}]\leq O(T^{-2})$. Note that
$X_{t}-\tau_{1}=\underbrace{\boldsymbol{\mu}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}}+\frac{1}{T_{1}}\sum_{s=1}^{T_{1}}\boldsymbol{y}_{s}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}}}_{:=A_{t}}+\underbrace{\frac{1}{T_{1}}\sum_{s=1}^{T_{1}}\boldsymbol{\mu}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{s}}_{:=\mathfrak{v}},$
so $\hat{\tau}_{1}-\tau_{1}=\frac{1}{T_{2}}\sum_{t=1}^{T_{2}}A_{t}+\mathfrak{v}$. We have $\mathbb{E}[\mathfrak{v}^{4}]\leq O(T^{-2})$, by Minkowski inequality, and
$\mathbb{E}[(\hat{\tau}_{1}-\tau_{1})^{4}]\leq 8\mathbb{E}[\mathfrak{v}^{4}]+8T_{2}^{-4}\mathbb{E}\left[\left(\sum_{t=1}^{T_{2}}A_{t}\right)^{4}\right]\leq 8T_{2}^{-4}\sum_{t_{1},t_{2}=1}^{T}\mathbb{E}[A_{t_{1}}^{2}A_{t_{2}}^{2}]+O(T^{-2}).$
Here, let’s prove that $\mathbb{E}[A_{t}^{4}]\leq O(1)$. For simplicity, let
$\mathfrak{y}_{T_{1}}:=\frac{1}{T_{1}}\sum_{s=1}^{T_{1}}\boldsymbol{y}_{s}.$
By Minkowski inequality, we have
$\mathbb{E}[A_{t}^{4}]\leq 8\mathbb{E}[(\boldsymbol{\mu}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}})^{4}]+8\mathbb{E}[(\mathfrak{y}_{T_{1}}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}})^{4}].$
Since $\mathbb{E}[(\boldsymbol{\mu}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}})^{4}]\leq O(1)$ by Assumption 2, it suffices to show that $\mathbb{E}[(\mathfrak{y}_{T_{1}}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}})^{4}]\leq O(1)$. Denote $\boldsymbol{B}:=\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Gamma}\mathfrak{y}_{T_{1}}\mathfrak{y}_{T_{1}}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Gamma}$, since $\mathfrak{y}_{T_{1}}$ and $\boldsymbol{y}_{t+T_{1}}$ are independent, then $\boldsymbol{B}$ and $\boldsymbol{y}_{t+T_{1}}$ are independent, by direct calculations, we have
$\mathbb{E}[(\mathfrak{y}_{T_{1}}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}})^{4}]=\mathbb{E}[(\boldsymbol{y}_{t+T_{1}}^{\prime}\boldsymbol{B}\boldsymbol{y}_{t+T_{1}})^{2}]$
$=\sum_{j_{1}\cdots j_{4}=1}^{m}\mathbb{E}[B_{j_{1},j_{2}}B_{j_{3},j_{4}}]\mathbb{E}[y_{j_{1},t+T_{1}}y_{j_{2},t+T_{1}}y_{j_{3},t+T_{1}}y_{j_{4},t+T_{1}}]$
$=\mathbb{E}[\operatorname{Tr}(\boldsymbol{B})^{2}]+2\mathbb{E}[\operatorname{Tr}(\boldsymbol{B}^{2})]+\sum_{j=1}^{m}\mathbb{E}[B_{j,j}^{2}](\mathbb{E}[y_{j;t+T_{1}}^{4}]-3),$
where
$\mathbb{E}[\operatorname{Tr}(\boldsymbol{B})^{2}]=\mathbb{E}[\operatorname{Tr}(\boldsymbol{B}^{2})]=\mathbb{E}[(\mathfrak{y}_{T_{1}}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\Gamma}\mathfrak{y}_{T_{1}})^{2}]$ (18)
$=T_{1}^{-2}\operatorname{Tr}(\boldsymbol{\Sigma}^{2})^{2}+2T_{1}^{-2}\operatorname{Tr}(\boldsymbol{\Sigma}^{4})+\sum_{j=1}^{m}(\Gamma_{\cdot j}^{\prime}\boldsymbol{\Sigma}\Gamma_{\cdot j})^{2}(\mathbb{E}[\mathfrak{y}_{j,T_{1}}^{4}]-3T_{1}^{-2}).$
According to Assumption 2, $\mathbb{E}[y_{j;t+T_{1}}^{4}]-3\leq|\kappa_{4}|<\infty$, then $\big|\mathbb{E}[\mathfrak{y}_{j,T_{1}}^{4}]-3T_{1}^{-2}\big|\leq|\kappa_{4}|T_{1}^{-2}$. This implies that $\mathbb{E}[\operatorname{Tr}(\boldsymbol{B})^{2}]=\mathbb{E}[\operatorname{Tr}(\boldsymbol{B}^{2})]\leq O(1)$. Similarly, since $\mathbb{E}[B_{j,j}^{2}]=\mathbb{E}[(\Gamma_{\cdot j}^{\prime}\boldsymbol{\Gamma}\mathfrak{y}_{T_{1}})^{4}]\leq O(T_{1}^{-2})$, then $\big|\sum_{j=1}^{m}\mathbb{E}[B_{j,j}^{2}](\mathbb{E}[y_{j;t+T_{1}}^{4}]-3)\big|\leq O(T^{-1})$ and $\mathbb{E}[A_{t}^{4}]\leq O(1)$. Consequently, we have
$\mathbb{E}[(\hat{\tau}_{1}-\tau_{1})^{4}]\leq 8T_{2}^{-4}\sum_{t_{1},t_{2}=1}^{T}\mathbb{E}[A_{t_{1}}^{2}A_{t_{2}}^{2}]+O(T^{-2})\leq O(T^{-2}).$
Next, let’s prove $\mathbb{E}[(\hat{\sigma}_{1}^{2}-\sigma_{1}^{2})^{2}]\leq O(T^{-1})$. Note that
$\hat{\sigma}_{1}^{2}-\sigma_{1}^{2}=\frac{1}{T_{2}}\sum_{t=1}^{T_{2}}(X_{t}-\tau_{1})^{2}-\sigma_{1}^{2}+\tau_{1}^{2}-\hat{\tau}_{1}^{2},$
where $\mathbb{E}[(\tau_{1}^{2}-\hat{\tau}_{1}^{2})^{2}]\leq\mathbb{E}[(\hat{\tau}_{1}-\tau_{1})^{4}]^{1/2}\mathbb{E}[(\hat{\tau}_{1}+\tau_{1})^{4}]^{1/2}\leq O(T^{-1})$, so it suffices to show that
$\frac{1}{T_{2}^{2}}\mathbb{E}\left[\left(\sum_{t=1}^{T_{2}}(X_{t}-\tau_{1})^{2}-\sigma_{1}^{2}\right)^{2}\right]=\frac{1}{T_{2}^{2}}\sum_{t_{1},t_{2}=1}^{T_{2}}\mathbb{E}[(X_{t_{1}}-\tau_{1})^{2}(X_{t_{2}}-\tau_{1})^{2}]-\sigma_{1}^{4}\leq O(T^{-1/2}).$
Since $(X_{t}-\tau_{1})^{2}=A_{t}^{2}+2A_{t}\mathfrak{v}+\mathfrak{v}^{2}$ and $\mathbb{E}[A_{t}^{4}]\leq O(1)$, for $t_{1}\neq t_{2}$,
$\mathbb{E}[A_{t_{1}}\mathfrak{v}^{3}]=\mathbb{E}[A_{t_{2}}\mathfrak{v}^{3}]=\mathbb{E}[A_{t_{1}}^{2}A_{t_{2}}\mathfrak{v}]=\mathbb{E}[A_{t_{1}}A_{t_{2}}^{2}\mathfrak{v}]=\mathbb{E}[A_{t_{1}}A_{t_{2}}\mathfrak{v}^{2}]=0.$
By Cauchy–Schwarz inequality, we have $\mathbb{E}[A_{t_{1}}^{2}\mathfrak{v}^{2}]\leq\mathbb{E}[A_{t_{1}}^{2}]^{1/2}\mathbb{E}[\mathfrak{v}^{2}]^{1/2}\leq O(T^{-1})$. Similarly, $\mathbb{E}[A_{t_{2}}^{2}\mathfrak{v}^{2}]\leq O(T^{-1})$. Therefore,
$\frac{1}{T_{2}^{2}}\sum_{t_{1}\neq t_{2}}^{T_{2}}\mathbb{E}[(X_{t_{1}}-\tau_{1})^{2}(X_{t_{2}}-\tau_{1})^{2}]=\frac{1}{T_{2}^{2}}\sum_{t_{1}\neq t_{2}}^{T_{2}}\mathbb{E}[A_{t_{1}}^{2}A_{t_{2}}^{2}]+O(T^{-1}).$
Moreover, by direct calculations, we have
$\mathbb{E}[A_{t_{1}}^{2}A_{t_{2}}^{2}]=\sigma_{1}^{4}+O(T^{-1}),\quad t_{1}\neq t_{2}.$
Consequently,
$\frac{1}{T_{2}^{2}}\sum_{t_{1}\neq t_{2}}^{T_{2}}(\mathbb{E}[(X_{t_{1}}-\tau_{1})^{2}(X_{t_{2}}-\tau_{1})^{2}]-\sigma_{1}^{4})=\frac{1}{T_{2}^{2}}\sum_{t_{1}\neq t_{2}}^{T_{2}}(\mathbb{E}[A_{t_{1}}^{2}A_{t_{2}}^{2}]-\sigma_{1}^{4})+O(T^{-1})\leq O(T^{-1}).$
Finally, $\mathbb{E}[(X_{t}-\tau_{1})^{4}]\leq 8\mathbb{E}[A_{t}^{4}]+8\mathbb{E}[\mathfrak{v}^{4}]\leq O(1)$, so $T_{2}^{-2}\sum_{t=1}^{T_{2}}\mathbb{E}[(X_{t}-\tau_{1})^{4}]-\sigma_{1}^{4}\leq O(1)$, which completes the proof of $\mathbb{E}[(\hat{\sigma}_{1}^{2}-\sigma_{1}^{2})^{2}]\leq O(T^{-1})$. ∎
A.2 Proof of Theorem 2.1
To conclude $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\overset{d}{\sim}\mathcal{B}(-\kappa_{1,T_{2}})$, it suffices to show that for any deterministic $a>0$
$\lim_{n\to\infty}\big|\mathbb{P}\big(\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|\leq a\big)-\mathbb{P}\big(\big|\mathcal{B}(-\kappa_{1,T_{2}})\big|\leq a\big)\big|=0.$
Actually, note that
$\mathbb{P}\big(\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|\leq a\big)=\mathbb{E}\big[1\big\{\big|\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})\big|\leq a\big\}\big],$
where the function $1_{|y|\leq a}$ is not smooth. Here, let’s approximate $1_{|y|\leq a}$ by a smooth function as follows. Let $\varphi_{a,\epsilon}(y)=\Phi((a-y)/\epsilon)-\Phi(-(a+y)/\epsilon)$, where $\epsilon>0$ and $\Phi(\cdot)$ is the CDF of the standard normal distribution, we have
$\lim_{\epsilon\downarrow 0}\sup_{y\in\mathbb{R}}|\varphi_{a,\epsilon}(y)-1_{|y|\leq a}|=0.$ (19)
It is easy to see $\varphi_{a,\epsilon}(y)=\varphi_{a,\epsilon}(-y)$ and
$\varphi_{a,\epsilon}^{\prime}(y)=\frac{1}{\sqrt{2\pi}\epsilon}\big(e^{-\frac{(a+y)^{2}}{2\epsilon^{2}}}-e^{-\frac{(a-y)^{2}}{2\epsilon^{2}}}\big),$
which implies that $\operatorname{sign}(\varphi_{a,\epsilon}^{\prime}(y))=-\operatorname{sign}(y)$.
Next, for any $t\in[0,1]$ and $x,\alpha,\beta\in\mathbb{R}$, let’s consider the following stochastic differential equation (SDE):
$dY_{s}^{x,t}=\alpha\operatorname{sign}(Y_{s}^{x,t})ds+\beta dB_{s},\quad s\in[t,1],$ (20)
where $Y_{t}^{x,t}=x$ and $B_{s}$ is the standard Brownian motion. According to [15], the above SDE admits a unique solution, and the density of $Y_{s}^{x,t}$ is given as follows:
$p_{x,t}(w;s)$ $=\frac{1}{\sqrt{2\pi(s-t)}\beta}e^{-\frac{(w/\beta-x/\beta)^{2}-2\alpha(s-t)(|w/\beta|-|x/\beta|)/\beta+\alpha^{2}(s-t)^{2}/\beta^{2}}{2(s-t)}}$
$-\alpha e^{2\alpha|w|/\beta^{2}}\int_{|w/\beta|+|x/\beta|+\alpha(s-t)/\beta}^{\infty}\frac{1}{\sqrt{2\pi(s-t)}\beta^{2}}e^{-\frac{u^{2}}{2(s-t)}}du,$ (21)
then define
$f_{t}(x):=\int_{\mathbb{R}}\varphi_{a,\epsilon}(w)p_{x,t}(w;1)dw.$
Lemma A.2.
The function $f_{t}(x)$ has the following properties:
  1. 1.
    The first and second order derivatives of $f_{t}(x)$ with respect to $x$ are uniformly bounded for all $t\in(0,1)$ and $x\in\mathbb{R}$.
  2. 2.
    There exists a constant $L>0$, independent of $t\in(0,1)$, such that —f_t(x_1)-f_t(x_2)—L—x_1-x_2—, x_1,x_2R.
  3. 3.
    For any $x\in\mathbb{R}$, $f_{t}(x)=f_{t}(-x)$ and $\operatorname{sign}(f^{\prime}_{t}(x))=-\operatorname{sign}(x)$ for any $x\in\mathbb{R}$.
  4. 4.
    Markov property: for any $t\in(0,1)$ and $r\in(1-t,1)$, $f_{t}(x)=\mathbb{E}[f_{t+r}(Y_{t+r}^{t,x})]$ for any $x\in\mathbb{R}$.
Readers can refer to [5, Lemma 5.1] for proofs of the above lemma. Here, we provide the Proof of Theorem 2.1 as below.
Proof.
First, let’s define a sequence of $\sigma$-fields $\{\mathcal{F}_{t}:t=0,1,\cdots,T_{2}\}$ such that $\mathcal{F}_{t}$ is the sigma field generated by $\{\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{T_{1}+t}\}$ for $t=0,1,\cdots,T_{2}$. For simplicity, we rewrite $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ by $\mathcal{M}_{T_{2},T_{2}}$ and define
$\mathfrak{N}_{t,T_{2}}:=\frac{\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}}{\sqrt{\tau_{1}^{2}+\sigma_{1}^{2}}}\mathcal{M}_{t,T_{2}},\quad t=1,\cdots,T_{2}.$
Since $\hat{\tau}_{1}\overset{\mathbb{P}}{\longrightarrow}\tau_{1}$ and $\hat{\sigma}_{1}\overset{\mathbb{P}}{\longrightarrow}\sigma_{1}$ by Lemma A.1, combining with (19), it suffices to show that for any deterministic $a>0$ and sufficiently small $\epsilon>0$
$\lim_{T\to\infty}$ $\big|\mathbb{E}\big[\varphi_{a,\epsilon}\big(\mathfrak{N}_{T_{2},T_{2}}\big)\big]-\mathbb{E}\big[\varphi_{a,\epsilon}\big(\mathcal{B}(-\kappa_{1,T_{2}})\big)\big]\big|=0.$
Here, we simplify $\varphi_{a,\epsilon}(\cdot)$ by $\varphi(\cdot)$, let $\alpha=-\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})$ and $\beta=1$ in (20), then we have $\mathbb{E}\big[\varphi_{a,\epsilon}\big(\mathcal{B}(-\kappa_{1,T_{2}})\big)\big]=f_{0}(0)$, and the above limit becomes
$\lim_{T\to\infty}\big|\mathbb{E}[\varphi(\mathfrak{N}_{T_{2},T_{2}})]-f_{0}(0)\big|=0.$
Note that
$\mathbb{E}[\varphi(\mathfrak{N}_{T_{2},T_{2}})]-f_{0}(0)=\sum_{t=1}^{T_{2}}\mathbb{E}\big[f_{t/T_{2}}(\mathfrak{N}_{t,T_{2}})-f_{(t-1)/T_{2}}(\mathfrak{N}_{t-1,T_{2}})\big]$
$=\underbrace{\sum_{t=1}^{T_{2}}\mathbb{E}\big[f_{t/T_{2}}(\mathfrak{N}_{t,T_{2}})-f_{t/T_{2}}(\mathfrak{N}_{t-1,T_{2}})\big]}_{:=I_{1,T_{2}}}+\underbrace{\sum_{t=1}^{T_{2}}\mathbb{E}\big[f_{t/T_{2}}(\mathfrak{N}_{t-1,T_{2}})-f_{(t-1)/T_{2}}(\mathfrak{N}_{t-1,T_{2}})\big]}_{:=I_{2,T_{2}}}.$
Let us first show that
$\big|I_{1,T_{2}}-\mathbb{E}[L_{t}(\mathfrak{N}_{t-1,T_{2}})]\big|\leq O(T^{-1/3})\quad\text{and}\quad\big|I_{2,T_{2}}+\mathbb{E}[L_{t}(\mathfrak{N}_{t-1,T_{2}})]\big|\leq O(T^{-1/2}),$
where
$L_{t}(\mathfrak{N}_{t-1,T_{2}}):=\frac{\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})}{T_{2}}\mathbb{E}[|f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})|]+\frac{1}{2T_{2}}\mathbb{E}[f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})].$
Note that
$\mathfrak{N}_{t,T_{2}}=\mathfrak{N}_{t-1,T_{2}}+\frac{\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}}{\sqrt{\tau_{1}^{2}+\sigma_{1}^{2}}}\frac{\theta_{t}X_{t}}{T_{2}}+\frac{\theta_{t}X_{t}}{\sqrt{T_{2}(\tau_{1}^{2}+\sigma_{1}^{2})}}:=\mathfrak{N}_{t-1,T}+\Delta_{t}.$
By Taylor expansion with Lagrange remainders, we have
$f_{t/T_{2}}(\mathfrak{N}_{t,T_{2}})$ $=f_{t/T_{2}}(\mathfrak{N}_{t-1,T_{2}})+f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\Delta_{t}+\frac{1}{2}f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})\Delta_{t}^{2}$
$+\frac{1}{2}\big(f_{t/T_{2}}^{\prime\prime}(\xi_{t})-f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})\big)\Delta_{t}^{2},$
where $\xi_{t}\in[\min\{\mathfrak{N}_{t-1,T_{2}},\mathfrak{N}_{t,T_{2}}\},\max\{\mathfrak{N}_{t-1,T_{2}},\mathfrak{N}_{t,T_{2}}\}]$.
Remainders: By Lemma A.2, there exists two constants $L_{2},L_{3}>0$ such that
$\sup_{y}|f_{t/T_{2}}^{\prime\prime}(y)|\leq L_{2}\quad\text{and}\quad\sup_{y_{1}\neq y_{2}}\frac{|f_{t/T_{2}}^{\prime\prime}(y_{1})-f_{t/T_{2}}^{\prime\prime}(y_{2})|}{|y_{1}-y_{2}|}\leq L_{3}.$
Since $|\xi_{t}-\mathfrak{N}_{t-1,T_{2}}|\leq|\Delta_{t}|$, then for $\eta=T^{-1/3}$,
$\big|f_{t/T_{2}}^{\prime\prime}(\xi_{t})-f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})\big|\leq\eta L_{3}1_{|\Delta_{t}|\leq\eta}+2L_{2}1_{|\Delta_{t}|>\eta}.$
Then
$\mathbb{E}\big[\big|f_{t/T_{2}}^{\prime\prime}(\xi_{t})-f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})\big|\Delta_{t}^{2}\big]\leq\eta L_{3}\mathbb{E}[|\Delta_{t}|^{2}1_{|\Delta_{t}|\leq\eta}]+2L_{2}\mathbb{E}[|\Delta_{t}|^{2}1_{|\Delta_{t}|>\eta}]$
$\leq 2\mathbb{E}[(T_{2}^{-2}(\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2})X_{t}^{2}/(\tau_{1}^{2}+\sigma_{1}^{2})+T_{2}^{-1}X_{t}^{2}/(\tau_{1}^{2}+\sigma_{1}^{2}))(\eta L_{3}1_{|\Delta_{t}|\leq\eta}+2L_{2}1_{|\Delta_{t}|>\eta})].$
According to Lemma A.1, we have $\mathbb{E}[X_{t}^{4}]\leq O(1)$ and $\mathbb{E}[\hat{\tau}_{1}^{4}],\mathbb{E}[\hat{\sigma}_{1}^{4}]\leq O(1)$, then by Cauchy–Schwarz inequality,
$\mathbb{E}[(\eta L_{3}1_{|\Delta_{t}|\leq\eta}+2L_{2}1_{|\Delta_{t}|>\eta})T_{2}^{-2}(\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2})X_{t}^{2}/(\tau_{1}^{2}+\sigma_{1}^{2})]\leq O(T^{-2}).$
Moreover, by Chebyshev inequality,
$T_{2}^{-1}\mathbb{E}[X_{t}^{2}1_{|\Delta_{t}|>\eta}]\leq T_{2}^{-1}\mathbb{E}[X_{t}^{4}]^{1/2}\mathbb{P}(|\Delta_{t}|>\eta)\leq O(T_{2}^{-1}\eta^{-2}\mathbb{E}[\Delta_{t}^{2}])\leq O(\eta^{-2}T^{-2})\leq O(T^{-4/3}).$
Combining with
$\eta T_{2}^{-1}\mathbb{E}[X_{t}^{2}1_{|\Delta_{t}|\leq\eta}]\leq O(\eta T^{-1}\mathbb{E}[X_{t}^{2}])\leq O(T^{-4/3}),$
we conclude that
$\mathbb{E}\big[\big|f_{t/T_{2}}^{\prime\prime}(\xi_{t})-f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})\big|\Delta_{t}^{2}\big]\leq O(T^{-4/3}).$ (22)
1st derivative: According to Lemma A.1, we have
$\mathbb{E}\big[\big(\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}/\sqrt{\tau_{1}^{2}+\sigma_{1}^{2}}-1\big)^{2}\big]$
$\leq\mathbb{E}\big[\big(\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}/\sqrt{\tau_{1}^{2}+\sigma_{1}^{2}}-1\big)^{2}\big(\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}/\sqrt{\tau_{1}^{2}+\sigma_{1}^{2}}+1\big)^{2}\big]$
$\leq\mathbb{E}\big[\big((\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2})/(\tau_{1}^{2}+\sigma_{1}^{2})-1\big)^{2}\big]\leq O(\mathbb{E}[(\hat{\tau}_{1}^{2}-\tau_{1}^{2})^{2}]+\mathbb{E}[(\hat{\sigma}_{1}^{2}-\sigma_{1}^{2})^{2}])\leq O(T^{-1}).$
Then
$\big|\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}X_{t}\sqrt{\hat{\tau}_{1}^{2}+\hat{\sigma}_{1}^{2}}/\sqrt{\tau_{1}^{2}+\sigma_{1}^{2}}]-\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}X_{t}]\big|$
$\leq O(T^{-1/2}\mathbb{E}[|f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})X_{t}|^{2}]^{1/2}).$
By Lemma A.2, $\sup_{y\in\mathbb{R}}|f_{t/T_{2}}^{\prime}(y)|\leq O(1)$, then $\mathbb{E}[|f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})X_{t}|^{2}]\leq O(\mathbb{E}[X_{t}^{2}])\leq O(1)$ and
$\big|\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}X_{t}\hat{\sigma}_{1}/\sigma_{1}]-\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}X_{t}]\big|\leq O(T^{-1/2}).$
Next, note that $\mathfrak{N}_{t-1,T_{2}}$ and $\theta_{t}$ are $\mathcal{F}_{t-1}$-measurable, then
$\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}X_{t}]=\mathbb{E}[f_{t/T}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}\mathbb{E}[X_{t}|\mathcal{F}_{t-1}]],$
and
$\mathbb{E}[X_{t}|\mathcal{F}_{t-1}]$ $=\frac{1}{T_{1}}\sum_{s=1}^{T_{1}}(\boldsymbol{x}_{s}^{\prime}\mathbb{E}[\boldsymbol{x}_{t+T_{1}}|\mathcal{F}_{t-1}]-d_{0}^{2})=\frac{1}{T_{1}}\sum_{s=1}^{T_{1}}(\boldsymbol{x}_{s}^{\prime}\boldsymbol{\mu}-d_{0}^{2}),$
where we use the fact that $\boldsymbol{x}_{t+T_{1}}$ is independent of $\mathcal{F}_{t-1}$, so $\operatorname{Var}(\mathbb{E}[X_{t}|\mathcal{F}_{t-1}])\leq O(T^{-1})$. By Lemma A.2, $\sup_{y\in\mathbb{R}}|f_{t/T_{2}}^{\prime}(y)|\leq O(1)$, it implies that $\operatorname{Var}(f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t})\leq O(1)$ and
$\big|\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}X_{t}]-\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}]\mathbb{E}[X_{t}]\big|\leq O(T^{-1/2}).$
Since $\operatorname{sign}(\varphi_{a,\epsilon}^{\prime}(y))=-\operatorname{sign}(y)$ and $\operatorname{sign}(\mathfrak{N}_{t-1,T_{2}})=\operatorname{sign}(\mathcal{M}_{t-1,T_{2}})$, by Lemma A.2, we have
$\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}]=\mathbb{E}[|f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})|],$
thus
$\big|\mathbb{E}[f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})\theta_{t}X_{t}]-\tau_{1}\mathbb{E}[|f_{t/T_{2}}^{\prime}(\mathfrak{N}_{t-1,T_{2}})|]\big|\leq O(T^{-1/2}),$
which implies that
$\mathbb{E}[f_{t/T}^{\prime}(\mathfrak{N}_{t-1,T})\Delta_{t}]=T_{2}^{-1}\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})\mathbb{E}[|f_{t/T}^{\prime}(\mathfrak{N}_{t-1,T})|]+O(T^{-3/2}).$ (23)
2nd derivative: Note that
$\mathbb{E}[X_{t}^{2}|\mathcal{F}_{t-1}]=\frac{1}{T_{1}^{2}}\sum_{s_{1},s_{2}=1}^{T_{1}}\mathbb{E}[\boldsymbol{x}_{s_{1}}^{\prime}\boldsymbol{x}_{t+T_{1}}\boldsymbol{x}_{t+T_{1}}^{\prime}\boldsymbol{x}_{s_{2}}|\mathcal{F}_{t-1}]-\frac{2d_{0}^{2}}{T_{1}}\sum_{s=1}^{T_{1}}\boldsymbol{x}_{s}^{\prime}\boldsymbol{\mu}+d_{0}^{4},$
where
$\mathbb{E}[\boldsymbol{x}_{s_{1}}^{\prime}\boldsymbol{x}_{t+T_{1}}\boldsymbol{x}_{t+T_{1}}^{\prime}\boldsymbol{x}_{s_{2}}|\mathcal{F}_{t-1}]=\boldsymbol{x}_{s_{1}}^{\prime}\boldsymbol{\mu}\boldsymbol{\mu}^{\prime}\boldsymbol{x}_{s_{2}}+\mathbb{E}[\boldsymbol{x}_{s_{1}}^{\prime}\boldsymbol{\Gamma}\boldsymbol{y}_{t+T_{1}}\boldsymbol{y}_{t+T_{1}}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{x}_{s_{2}}|\mathcal{F}_{t-1}]$
$=\boldsymbol{x}_{s_{1}}^{\prime}\boldsymbol{\mu}\boldsymbol{\mu}^{\prime}\boldsymbol{x}_{s_{2}}+\boldsymbol{x}_{s_{1}}^{\prime}\boldsymbol{\Sigma}\boldsymbol{x}_{s_{2}}=\|\boldsymbol{\mu}\|_{2}^{4}+\|\boldsymbol{\mu}\|_{2}^{2}\boldsymbol{\mu}^{\prime}\boldsymbol{\Gamma}(\boldsymbol{y}_{s_{1}}+\boldsymbol{y}_{s_{2}})+\boldsymbol{y}_{s_{1}}^{\prime}\boldsymbol{\mu}\boldsymbol{\mu}^{\prime}\boldsymbol{y}_{s_{2}}$
$+\boldsymbol{\mu}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\mu}+\boldsymbol{\mu}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\Gamma}(\boldsymbol{y}_{s_{1}}+\boldsymbol{y}_{s_{2}})+\boldsymbol{y}_{s_{1}}^{\prime}\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\Gamma}\boldsymbol{y}_{s_{2}}.$
Thus, we obtain that
$\mathbb{E}[X_{t}^{2}|\mathcal{F}_{t-1}]=\tau_{1}^{2}+\boldsymbol{\mu}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\mu}+2(\tau_{1}\boldsymbol{\mu}^{\prime}+\boldsymbol{\mu}^{\prime}\boldsymbol{\Sigma})\boldsymbol{\Gamma}\mathfrak{y}_{T_{1}}+\mathfrak{y}_{T_{1}}^{\prime}(\boldsymbol{\mu}\boldsymbol{\mu}^{\prime}+\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\Gamma})\mathfrak{y}_{T_{1}},$
where $\mathfrak{y}_{T_{1}}=T^{-1}\sum_{s=1}^{T_{1}}\boldsymbol{y}_{s}$. According to (18), we know that
$\operatorname{Var}\big(\mathfrak{y}_{T_{1}}^{\prime}(\boldsymbol{\mu}\boldsymbol{\mu}^{\prime}+\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\Gamma})\mathfrak{y}_{T_{1}}\big)$
$\leq\frac{2\operatorname{Tr}((\boldsymbol{\mu}\boldsymbol{\mu}^{\prime}+\boldsymbol{\Gamma}^{\prime}\boldsymbol{\Sigma}\boldsymbol{\Gamma})^{2})}{T_{1}^{2}}+\sum_{j=1}^{m}(\mu_{j}^{2}+\Gamma_{\cdot j}^{\prime}\boldsymbol{\Sigma}\Gamma_{\cdot j})^{2}(\mathbb{E}[\mathfrak{y}_{j,T_{1}}^{4}]-3T_{1}^{-2})\leq O(T^{-1}).$
Combining with the fact that $\sup_{y\in\mathbb{R}}|f_{t/T_{2}}^{\prime\prime}(y)|\leq O(1)$ by Lemma A.2, we derive that
$\big|\mathbb{E}[f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})X_{t}^{2}]-\mathbb{E}[X_{t}^{2}]\mathbb{E}[f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})]\big|\leq O(T^{-1/2}),$
i.e.
$\mathbb{E}[f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})X_{t}^{2}]=(\tau_{1}^{2}+\sigma_{1}^{2})\mathbb{E}[f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})]+O(T^{-1/2}).$
Consequently, we can further obtain that
$\mathbb{E}[f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})\Delta_{t}^{2}]=T_{2}^{-1}\mathbb{E}[f_{t/T_{2}}^{\prime\prime}(\mathfrak{N}_{t-1,T_{2}})]+O(T^{-3/2}).$ (24)
In summary, combining (22), (23) and (24), we conclude that that
$\Bigg|I_{1,T_{2}}-\sum_{t=1}^{T_{2}}L_{t}(\mathfrak{N}_{t-1,T})\Bigg|$
$\leq\sum_{t=1}^{T_{2}}\big|\mathbb{E}\big[f_{t/T_{2}}(\mathfrak{N}_{t,T_{2}})-f_{t/T_{2}}(\mathfrak{N}_{t-1,T_{2}})\big]-L_{t}(\mathfrak{N}_{t-1,T_{2}})\big|\leq O(T^{-1/3}).$
On the other hand, by Lemma A.2, for any $x\in\mathbb{R}$,
$f_{(t-1)/T_{2}}(x)=\mathbb{E}[f_{t/T_{2}}(Y_{t/T_{2}}^{x,(t-1)/T_{2}})].$
By $\mathrm{It\hat{o}}$’s formula and Dynkin’s formula, we have
$f_{(t-1)/T_{2}}(x)-f_{t/T_{2}}(x)=\mathbb{E}\big[f_{t/T_{2}}(Y_{t/T_{2}}^{x,(t-1)/T_{2}})\big]-f_{t/T_{2}}(x)$
$=\mathbb{E}\Bigg[\int_{(t-1)/T_{2}}^{t/T_{2}}-\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})\operatorname{sign}(Y_{s}^{x,(t-1)/T_{2}})f_{t/T_{2}}^{\prime}(Y_{s}^{x,(t-1)/T_{2}})$
$+\frac{1}{2}f_{t/T_{2}}^{\prime\prime}(Y_{s}^{x,(t-1)/T_{2}})ds+\int_{(t-1)/T_{2}}^{t/T_{2}}f_{t/T_{2}}^{\prime}(Y_{s}^{x,(t-1)/T_{2}})dB_{s}\Bigg]$
$=\mathbb{E}\Bigg[\int_{(t-1)/T_{2}}^{t/T_{2}}\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})\big|f_{t/T_{2}}^{\prime}(Y_{s}^{x,(t-1)/T_{2}})\big|+\frac{1}{2}f_{t/T_{2}}^{\prime\prime}(Y_{s}^{x,(t-1)/T_{2}})ds\Bigg].$
Thus, by Lemma A.2, we conclude that
$\sum_{t=1}^{T_{2}}\sup_{x\in\mathbb{R}}\big|f_{(t-1)/T_{2}}(x)-f_{t/T_{2}}(x)-L_{t}(x)\big|$
$\leq\tau_{1}\sum_{t=1}^{T}\sup_{x\in\mathbb{R}}\mathbb{E}\Bigg[\int_{(t-1)/T_{2}}^{t/T_{2}}\big|f_{t/T_{2}}^{\prime}(Y_{s}^{x,(t-1)/T_{2}})-f^{\prime}_{t/T_{2}}(x)\big|ds\Bigg]$
$+\frac{1}{2}\sum_{t=1}^{T_{2}}\sup_{x\in\mathbb{R}}\mathbb{E}\Bigg[\int_{(t-1)/T_{2}}^{t/T_{2}}\big|f_{t/T_{2}}^{\prime\prime}(Y_{s}^{x,(t-1)/T_{2}})-f_{t/T_{2}}^{\prime\prime}(x)\big|ds\Bigg]$
$\leq O\Bigg(\frac{1}{T_{2}}\sum_{t=1}^{T_{2}}\sup_{x\in\mathbb{R}}\mathbb{E}\Bigg[\sup_{s\in[(t-1)/T_{2},t/T_{2}]}\big|Y_{s}^{x,(t-1)/T_{2}}-x\big|\Bigg]\Bigg)$
$\leq O\Bigg(\frac{1}{T_{2}}\sum_{t=1}^{T_{2}}\mathbb{E}\Bigg[\frac{\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})}{T_{2}}+\sup_{s\in[(t-1)/T_{2},t/T_{2}]}\big|B_{s}-B_{(t-1)/T_{2}}\big|\Bigg]\Bigg)\leq O(T^{-1/2}),$
which implies that $\big|I_{2,T_{2}}+\sum_{t=1}^{T_{2}}L_{t}(\mathfrak{N}_{t-1,T_{2}})\big|\leq O(T^{-1/2})$, so $\lim_{T\to\infty}|I_{1,T_{2}}-I_{2,T_{2}}|=0$. Finally, according to (21), let $\alpha=\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})$ and $\beta=1$, for any $a\geq 0$, we have
$\int_{-a}^{a}p_{0,0}(w;1)dw=\Phi(\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})+a)$
$-e^{-2a\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})}\Phi(\tau_{1}(1+\sqrt{T_{2}/(\tau_{1}^{2}+\sigma_{1}^{2})})-a),$
which completes the proof. ∎